OpenStudy (anonymous):

Solve the equation: xy=6 x^2+y^2=13

OpenStudy (anonymous):

solve first equation for y y=6/x plug that expression in for y in the second equation x^2+(6/x)^2=13 now can you try and solve for x?

OpenStudy (jdoe0001):

\(\bf x^2+\left(\cfrac{6}{x}\right)^2=13x^2\implies x^2+\cfrac{36}{x^2}=13\\ \quad \\ \textit{multiplying both sides by }x^2\\ \quad \\ x^2\left(x^2+\cfrac{36}{x^2}\right)=x^2(13)\\ \quad \\\implies x^4+36=13x^2\implies [x^2]^2-13[x]^2+36=0\) factor the quadratic to get the 4 values for "x"

OpenStudy (jdoe0001):

hmm added a premature \(x^2 \).

OpenStudy (jdoe0001):

\(\bf x^2+\left(\cfrac{6}{x}\right)^2=13\implies x^2+\cfrac{36}{x^2}=13\\ \quad \\ \textit{multiplying both sides by }x^2\\ \quad \\ x^2\left(x^2+\cfrac{36}{x^2}\right)=x^2(13)\\ \quad \\\implies x^4+36=13x^2\implies [x^2]^2-13[x]^2+36=0\)

OpenStudy (jdoe0001):

factor it to get the values